# ARTIST: Improving the Generation of Text-rich Images with Disentangled Diffusion Models

### v2

@@ 4 Experiments
4.1 Implementation Details
Our experiments are conducted on 8 Nvidia A100 GPUs, with Hugging Face Diffusers [39]. Both text and visual modules are initialized from the pretrained Stable Diffusion [29] checkpoint. The text module is first pretrained on our word-level synthetic dataset for 400,000 steps, then further finetuned for 200,000 steps on our sentence-level synthetic dataset. With the frozen text module, our visual module is trained for 250,000 steps on MARIO-10M dataset [6]. AdamW [18] optimizer is used in all training stages, with a learning rate of 1e-5 and a weight decay of 1e-2. The batch size is set to 128.
We first compare our proposed framework with previous methods including Stable Diffusion [29] (denoted as SD), ControlNet [49], DeepFloyd [35] and TextDiffuser [6] in terms of common metrics such as OCR accuracy and FID. We also finetune a Stable Diffusion on MARIO-10M dataset for a more comprehensive comparison. Then, we directly compare our method with the latest approaches like AnyText [38] and TextDiffuser 2 [5] through human evaluation, as it encapsulates the capabilities of the above traditional metrics while also assessing subjective aspects crucial for user experience and practical application effectiveness. Two variants of our proposed framework are evaluated, which are denoted as ARTIST-TD and ARTIST, respectively, indicating whether LLM is utilized. Specifically, ARTIST-TD directly uses the pretrained transformer from TextDiffuser instead of LLM to obtain bounding boxes and keywords based on input prompts. Thus, comparing ARTIST-TD with TextDiffuser can straightforwardly show the effectiveness of our training strategy, as they share exactly the same layout and keyword conditions.
Note that although our framework has two separate diffusion models, our computation requirement is still similar to the previous SOTA TextDiffuser. This is because TextDiffuser also utilizes an extra U-Net, which is designed for character-aware loss, as a regularization term.
@@ 4.1 Implementation Details
Metrics SD finetuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@ 4.2 Main Results
To start with, we conduct experiments on the MARIO-Eval benchmark proposed in  [6], which contains 5,414 prompts in total. 4 images are generated for each prompt to compute the CLIP score [24], Fréchet Inception Distance (FID) [9] and OCR evaluations. Specifically, CLIP score is obtained by calculating the cosine similarity between generated images and prompts by using the features extracted with the pretrained ViT/B-32 CLIP model. OCR evaluation is performed with MaskSpotterv3 [16] following [6]. The main reuslts are presented in Table 1, from which we can see that our ARTIST outperforms the previous SOTA TextDiffuser in all metrics, even without the help of LLM. Some generated examples are provided in Figure 1, more results are provided in the Appendix because of the limited space.
@@
As the reader may notice from Figure 1 and 4, most keywords in MARIO-Eval prompts are enclosed by quotation marks. In other words, a model can easily obtain promising accuracy in keyword prediction even if it is over-fitting and simply extracts words enclosed by quotation marks. Thus MARIO-Eval benchmark is not a reasonable benchmark to justify model performance on open-domain instructions. To this end, we propose a new ARTIST-Eval benchmark, which contains 500 pairs of prompts and keywords. Details about constructing the benchmark and some examples are provided in the Appendix B. The keywords in the ARTIST-Eval benchmark, which are expected to be shown in the resulting images, are designed to be contained in the prompts so that we can compare all the methods fairly. However, keywords are not always enclosed by quotation marks. Similar to the MARIO-Eval benchmark, 4 images are generated for each prompt to calculate the CLIP score and OCR evaluations. The results on the ARTIST-Eval benchmark are presented in Table 2. From the results, we can see that both ARTIST-TD and ARTIST outperform related methods, while ARTIST obtains a huge performance boost, indicating the benefits of using LLM.
@@
Metrics SD finetuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@
Figure 5: Generated examples in inpainting task, where the masked regions are indicated by red rectangles. Prompts used in these examples are “a book cover of Bed Times", “a book cover for Kansas State", “a poster for A Good Princess" and “a poster for The Man Who Said YES".
@@
To quantitatively evaluate the contribution of LLM, we conduct experiments on keyword identification. Different models are asked to extract keywords from prompts of the ARTIST-Eval benchmark. The results are shown in Table 3, where we can conclude that LLM indeed leads to huge improvement in detecting keywords.
@@
Accuracy ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ ) F-measure ()($\uparrow$)( ↑ )
@@
Similar to TextDiffuser, the proposed ARTIST can also be trained to perform an image-inpainting task. We trained another ARTIST following the setting from Chen et al. [6] so that the model is asked to perform image inpainting instead of image generation with a probability of 0.5 during training. The resulting ARTIST can perform both tasks at inference time. Some generated examples are provided in Figure 5, along with the comparison between ARTIST and TextDiffuser. Similarly to the whole-image generation task, our ARTIST leads to more accurately rendered texts and more harmonized images.
@@ 4.4 Ablation Studies
Although a character segmentation mask is used in our implementation following  [6], the proposed ARTIST should lead to improvements no matter whether the character segmentation mask is used or not. To verify this claim, we conduct an ablation study to show the robustness of ARTIST. Specifically, we trained an ARTIST model without the character-level segmentation mask, which is to be compared with the finetuned Stable Diffusion model. For all the models, we apply 50 sampling steps and classifier-free guidance [11] of $\omega\in\{2.5,5,7.5,10,15\}$ during inference. In general, a larger $\omega$ leads to better FID and CLIP score, while it may lead to a worse OCR evaluation. The comparison is provided in Figure 6, from which we can conclude that our ARTIST leads to improvement across different settings because it leads to better alignment of the image and the text, image fidelity, and OCR accuracy compared to baselines.
@@
Features from FID ()($\downarrow$)( ↓ ) CLIP Score ()($\uparrow$)( ↑ ) OCR Evaluation
Accuracy ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ ) F-measure ()($\uparrow$)( ↑ )
U-Net Encoder 37.454 0.3455 0.4935 0.7312 0.7293 0.7302
U-Net Decoder 36.579 0.3466 0.6298 0.8237 0.7986 0.8110
@@ 5 Conclusions
We proposed ARTIST, a novel framework that significantly enhances the text-rendering ability of diffusion models. Our proposed framework utilizes pretrained large language models to infer the user’s intention, provide accurate prompts, and improve the interactive experience. We also introduced a disentangled architecture design and training strategy, leading to better learning of text structure and visual appearance. Our experimental results demonstrate that ARTIST outperforms the previous state-of-the-art in terms of image fidelity, image-prompt alignment, and accuracy of generated texts. In the future, we aim to further improve the disentangled learning, investigate the interpretability of the disentangled representations learned by ARTIST and their potential for downstream tasks. Overall, we believe that ARTIST represents a significant step forward in the field of text-rich image generation and has the potential to enable new applications in the future.
Appendix A Limitation and Broader Impact
This paper presents work whose goal is to enhance the capabilities of generative models in producing images with legible and contextually appropriate text. While this advancement primarily aims to contribute to the field of Machine Learning, it also holds potential for societal impact. For example, it could streamline graphic design processes, improve accessibility through clearer visual communication, and foster creative expression across digital platforms. However, we acknowledge that the ability to generate realistic text within images could be misused to create misleading information or deepfakes. Therefore, we encourage the development of robust detection methods alongside such generative technologies and support the creation of ethical guidelines governing their use.
Appendix B ARTIST-Eval benchmark
@@
Appendix C More Results
@@
Appendix D Examples from the survey

### v3

@@ 4 Experiments
Metrics SD Fine-tuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@
4.1 Implementation Details
Our experiments are conducted on 8 Nvidia A100 GPUs, with Hugging Face Diffusers [46]. Both text and visual modules are initialized from the pre-trained Stable Diffusion [34] checkpoint. The text module is first pre-trained on our word-level synthetic dataset for 400,000 steps, then further fine-tuned for 200,000 steps on our sentence-level synthetic dataset. With the frozen text module, our visual module is trained for 250,000 steps on MARIO-10M dataset [7]. AdamW [24] optimizer is used in all training stages, with a learning rate of 1e-5 and a weight decay of 1e-2. The batch size is set to 128.
We first compare our proposed framework with previous methods, including Stable Diffusion [34] (denoted as SD), ControlNet [70], DeepFloyd [41] and TextDiffuser [7] in terms of common metrics such as OCR accuracy and FID. We also finetune a Stable Diffusion on MARIO-10M dataset for a more comprehensive comparison. Then, we directly compare our method with the latest approaches like AnyText [45] and TextDiffuser 2 [6] through human evaluation, as it encapsulates the capabilities of the above traditional metrics while also assessing subjective aspects crucial for user experience and practical application effectiveness. Two variants of our proposed framework are evaluated, which are denoted as ARTIST-TD and ARTIST, respectively, indicating whether LLM is utilized. Specifically, ARTIST-TD directly uses the pretrained transformer from TextDiffuser instead of LLM to obtain bounding boxes and keywords based on input prompts. Thus, comparing ARTIST-TD with TextDiffuser can straightforwardly show the effectiveness of our training strategy, as they share exactly the same layout and keyword conditions.
Note that although there are two separate diffusion models in our framework, our computation requirement is still similar to the previous SOTA TextDiffuser. This is because TextDiffuser also utilizes an extra U-Net, which is designed for character-aware loss as a regularization term.
@@ 4.2 Main Results
To start with, we conduct experiments on MARIO-Eval benchmark proposed in  [7], which contains 5,414 prompts in total. 4 images are generated for each prompt to compute the CLIP score [30], Fréchet Inception Distance (FID) [12] and OCR evaluations. Specifically, CLIP score is obtained by calculating the cosine similarity between generated images and prompts by using the features extracted with pre-trained ViT/B-32 CLIP model. OCR evaluation is performed with MaskSpotterv3 [22] following [7]. The main results are presented in Table 1, from which we can see that our ARTIST outperforms the previous SOTA TextDiffuser in all metrics, even without the help of LLM. Some generated examples are provided in Figure 1, and more results are provided in the Appendix because of the limited space.
@@
As the reader may notice from Figure 1 and 4, most keywords in MARIO-Eval prompts are enclosed by quotation marks. In other words, a model can easily obtain promising accuracy in keyword prediction even if it is over-fitting and simply extracts words enclosed by quotation marks. Thus MARIO-Eval benchmark is not a reasonable benchmark to justify model performance on open-domain instructions. To this end, we propose a new ARTIST-Eval benchmark, which contains 500 pairs of prompts and keywords. Details about constructing the benchmark and some examples are provided in the Appendix A. The keywords in ARTIST-Eval benchmark, which are expected to be shown in the resulting images, are designed to be contained in the prompts so that we can compare all the methods fairly. However, keywords are not always enclosed by quotation marks. Similar to MARIO-Eval benchmark, 4 images are generated for each prompt to calculate CLIP score and OCR evaluations. The results on ARTIST-Eval benchmark are presented in Table 2. From the results we can see that both ARTIST-TD and ARTIST outperform related methods, while ARTIST obtains a huge performance boost, indicating the benefits of using LLM.
@@
Metrics SD Fine-tuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@
Figure 5: Generated examples in inpainting task, where the masked regions are indicated by red rectangles. Prompts used are “a book cover of Bed Times", “a book cover for Kansas State", “a poster for A Good Princess" and “a poster for The Man Who Said YES".
@@
To quantitatively evaluate the contribution of LLM, we conduct experiments of keywords identification. Different models are asked to extract keywords from prompts of ARTIST-Eval benchmark. The results are shown in Table 3, where we can conclude that LLM indeed leads to huge improvement in detecting keywords.
@@
Acc ()($\uparrow$)( ↑ ) P ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) F1 ()($\uparrow$)( ↑ )
@@
Similar to TextDiffuser, the proposed ARTIST can also be trained to perform an image-inpainting task. We trained another ARTIST following the setting from Chen et al. [7] so that the model is asked to perform image inpainting instead of image generation with a probability of 0.5 during training. The resulting ARTIST is able to perform both tasks at inference time. Some generated examples are provided in Figure 5, along with the comparison between ARTIST and TextDiffuser. Similarly to the whole-image generation task, our ARTIST leads to more accurately rendered texts and more harmonized images.
U-Net Encoder U-Net Decoder
FID ()($\downarrow$)( ↓ ) 37.454 36.579
CLIP Score ()($\uparrow$)( ↑ ) 0.3455 0.3466
OCR Evaluation
Accuracy ()($\uparrow$)( ↑ ) 0.4935 0.6298
Precision ()($\uparrow$)( ↑ ) 0.7312 0.8237
Recall ()($\uparrow$)( ↑ ) 0.7293 0.7986
F-measure ()($\uparrow$)( ↑ ) 0.7302 0.8110
@@ 4.4 Ablation Studies
Although character segmentation mask is used in our implementation following  [7], the proposed ARTIST should lead to improvements no matter whether character segmentation mask is used or not. To verify this claim, we conduct an ablation study to show the robustness of ARTIST. Specifically, we trained an ARTIST model without character-level segmentation mask, which is to be compared with the fine-tuned Stable Diffusion model. For all the models, we apply 50 sampling steps and classifier-free guidance [14] of $\omega\in\{2.5,5,7.5,10,15\}$ during inference. In general, a larger $\omega$ leads to better FID and CLIP scores, while it may lead to a worse OCR evaluation. The comparison is provided in Figure 6, from which we can conclude that our ARTIST leads to improvement across different settings because it leads to better alignment of the image and the text, image fidelity, and OCR accuracy compared to baselines.
@@ 5 Conclusions
We proposed ARTIST, a novel framework that significantly enhances the text-rendering ability of diffusion models. Our proposed framework utilizes pretrained large language models to infer the user’s intention, provide accurate prompts, and improve the interactive experience. We also introduced a disentangled architecture design and training strategy which leads to better learning of both text structure and visual appearance. Our experimental results demonstrate that ARTIST outperforms the previous state-of-the-art in terms of image fidelity, image-prompt alignment, and accuracy of generated texts. In the future, we aim to improve the disentangled learning further, investigate the interpretability of the disentangled representations learned by ARTIST and their potential for downstream tasks. Overall, we believe that ARTIST represents a significant step forward in the field of text-rich image generation and has the potential to enable new applications in the future.
Appendix A ARTIST-Eval benchmark
@@
Appendix B More Results
@@
Appendix C Examples from the survey